The properties of a closed system change following the relation between pressure and volume as pV= 3.0, where P is in bar V is in m^3. Calculate the work done when the pressure increases from 1.5 bar to 7.5 bar.
1
Expert's answer
2020-11-13T10:51:12-0500

The relation between pressure(P) and volume (V) of a closed system is given as

"P\times V=3=constant"

"P=\frac{3}{V}" ---------------------------------------(1)

"P_1\times V_1=P_2 \times V_2"

when P=1.5 bar=1.5"\times 10^5 Pa" => V="\frac{3}{1.5\times 10^5}=2\times 10^{-5}" "m^3"

when P=7.5 bar=1.5"\times 10^5 Pa" => V="\frac{3}{7.5\times 10^5}=0.4\times 10^{-5} m^3" "m^3"

The work done during the process is given by

"W=\int_{V_1 }^{V_2}PdV=\int_{2\times 10^{-5} }^{0.5\times 10^{-5}}\frac{3}{V}]dV=3\times ln(\frac{0.4}{2})=-3\times 1.6 J=-4.828J"

Negative sign indicates work done on the system.